OpenStudy (anonymous):

Choose an equation for the line that is parallel to the given line and passes through the given point for questions 1 and 2. y = x – 10; (15, –16) y = –5x + 19 y = x + y = x – 19 y = 5x – 19 Choose an equation for the line that is parallel to the given line and passes through the given point for questions 1 and 2. y = x – 10; (15, –16) y = –5x + 19 y = x + y = x – 19 y = 5x – 19 @Mathematics

OpenStudy (anonymous):

y=1(x-15) -16 y=x-15-16 y=x-31